OpenStudy (anonymous):

The equation y = 55x + 85 describes the amount of money a plumber charges, where y is the total amount he charges and x is the number of hours he works. The table shows the total amount of money an electrician charges. How much does each charge per hour? Number of hours worked 1 2 4 7 Total amount charged 115 180 310 505 The plumber charges $ per hour. The electrician charges $ per hour.

OpenStudy (anonymous):

Does the plumber charge $55?

OpenStudy (anonymous):

and the electrician $85?

OpenStudy (welshfella):

for the plumber y = 55x + 85 for 1 hour the charge is 55(x) = 55(1) = 55

OpenStudy (anonymous):

does the electrician charge $60, $65 or $85?

OpenStudy (welshfella):

for the electrician y = ax + z 115 = a + z 180 = 2a + z where a is charge per hour and z = standard charegr subtract first equation for the second a = 180 - 115 = 65 electriciam charges 65 dollars per hour

OpenStudy (anonymous):

Okay! Thank You! ^.^

OpenStudy (welshfella):

yw